it is very important that you show your work from the calculations (right down what you used as PV, i FV, n, cash flows,.....etc. in your calculation)

When you purchased your house 4 years ago, you took out a $300,000, 30 year FRM loan at 7.5%. You also paid 2 points to get this loan. Now you want to sell your house and buy another one. In order to do so, you must pay off the existing one. The loan came with 1% prepayment penalty. What has been your effective interest rate on this loan?
Answer #1

Step 1: Monthly payment
PV=-300000
I/Y=7.5%/12
N=12*30
FV=0
CPT PMT=2097.643526

Step 2: Loan balance
N=12*4
I/Y=7.5%/12
PV=-300000
PMT=2097.643526
CPT FV=287581.864926

Step 3: Interest rate
PV=-300000*(1-2%)=-294000.00
FV=287581.864926*(1+1%)=290457.683575
PMT=2097.643526
N=12*4
CPT I/Y=0.692237%

Hence, effective interest rate=(1+0.692237%)^12-1=8.630523%